Solve by Gaussian elimination method the following system of equations :

x+y+z+t=5

x-y+z+t=1

x + z + t = 3


1
Expert's answer
2020-11-17T15:34:15-0500

make a matrix:

"\begin{pmatrix}
 1& 1&1&1&&5 \\
 1&-1&1&1&&1\\
1&0&1&1&&3
\end{pmatrix}"

make this matrix simple form:

line 3 minus line 2:

"\begin{pmatrix}
 1& 1&1&1&&5 \\
 1&-1&1&1&&1\\
0&1&0&0&&2
\end{pmatrix}"

line 2 minus line 1:

"\begin{pmatrix}
 1& 1&1&1&&5 \\
 0&-2&0&0&&-4\\
0&1&0&0&&2
\end{pmatrix}"

line 2 divided by -2:

"\begin{pmatrix}
 1& 1&1&1&&5 \\
 0&1&0&0&&2\\
0&1&0&0&&2
\end{pmatrix}"

lines 2 and 3 are equal, so we remove the third line:

"\begin{pmatrix}
 1& 1&1&1&&5 \\
 0&1&0&0&&2\\
\end{pmatrix}"

this matrix has a simple form, back to system equations:

"x+y+z+t=5,"

"y=2;"


"x=5-2-z-t,"

"y=2;"


z and t are free variables

"z=h1,"

"t=h2;"


solution to the system of equations:

"x=3- h1-h2,"

"y=2,"

"z=h1,"

"t=h2;"

(where h1 and h2 are any real numbers)